# Repurposing Pre-trained Video Diffusion Models for Event-based Video Interpolation

### v1

@@ Abstract
Figure 1: We compare our proposed approach RE-VDM, which adapts a pre-trained video diffusion model for event-based video frame interpolation on unseen real-world data, with three baselines: frame-only interpolation method RIFE [15], test-time optimization of a pre-trained video diffusion model via Time Reversal [10], and event-based interpolation method CBMNet-Large [18], trained on the same dataset as our method. The left-most column shows the start frame, end frame, and reference interpolation frames overlaid with events. Leveraging data priors in the pre-trained video diffusion model and the diffusion process, along with motion guidance controlled by input events, our approach demonstrates superior generalization performance on unseen real-world frames with substantial motion.
@@ 2.1 Event Camera and Video Frame Interpolation
Event cameras are neuromorphic sensors that capture the brightness changes in the scene, and the brightness changes are primarily due to the camera’s motion and object motion in the scene. For this reason, events captured by event cameras can be used to compute motion-related information, for example, optical flow, motion segmentation, and ego-motion [6, 4, 3, 5, 24, 25, 27, 45, 44]. Various methods have been proposed to reconstruct intensity videos from event streams alone, demonstrating that while feasible, the quality of such reconstructions is significantly limited by camera and object motion [9, 30]. Combining events with images extends beyond event-only applications, offering powerful capabilities for tasks like image deblurring and video frame interpolation [37, 38, 42, 36]. This approach leverages events as cues to bridge the gap between frames, thus avoiding the ill-posed challenges faced by frame-only methods [15, 28, 21, 17, 1, 19] when interpolating between frames with substantial motion.
@@ 2.2 Diffusion and Video Diffusion Models
Here, $\alpha(t),\sigma(t)$ represents a noise schedule. Diffusion models have demonstrated remarkable image generation quality and diversity. Stable Diffusion [31] shifts the generation process to a low-dimensional latent space, greatly reducing computational cost. Trained on massive online text-image pair datasets [33], Stable Diffusion exhibits strong image generation capabilities.
Video diffusion models [14, 8] add temporal layers on top of existing image diffusion models to jointly denoise multiple consistent image frames. Stable Diffusion Video (SVD) [8] is built on top of latent diffusion model and trained on massive video datasets, serving as a strong foundational model for video generation tasks.
@@ 2.3 Controllable Diffusion Generation Through Fine-Tuning
There have been numerous works on generation conditioned on additional user inputs or control [31, 32, 40]. Stable Diffusion [31] uses cross-attention layers in the diffusion U-Net to inject text control. Dreambooth [32] showed that a diffusion model can be fine-tuned for conditional control. ControlNet [40] modifies a trainable copy of the original diffusion U-Net for conditional control. In addition to fine-tuning diffusion models for added controllability, various training-free methods have been proposed to alter the sampling process of image and video diffusion models for more flexible and generalized content synthesis [2, 10, 22, 43]. Latent diffusion models have limitations on generating a limited set of resolutions in a downsampled latent space. MultiDiffusion [2] provides a solution for resolution-agnostic generation by tile-based denoising. When combined with conditional control, MultiDiffusion has shown to be effective in generating synthetic detailed images under the condition of fine-grained condition control [16]. On the video generation side,  [10] proposes a test-time sampling method that empowers video diffusion models to control camera and object motion.  [43] enables consistent long-frame generation outside of the trained frame length of a video diffusion model, through simultaneously generating multiple video sequences, and fusing intermediate denoising results of overlapping frames.
@@ 3.1 Pipeline Overview
Figure 2 presents an overview of our pipeline RE-VDM, for adapting pre-trained video diffusion models to event-based video generation and interpolation tasks. The key difference is that interpolation applies our EVDS on both the start frame $I_{s}$ for forward-time video denoising/generation and the end frame $I_{e}$ for backward-time video denosing/generation. At each denoising step, forward-backward consistency is enforced to achieve accurate interpolation. Details and design of each key component are explained in the following subsections.
@@ 3.3 Events as the Control Condition
To introduce events as the control condition for the video generation process, we incorporate two main components: 1) Event Representation and 2) Event Encoder. For event representation, we adopt a multi-stack approach, inspired by [26], which captures both fast- and slow-moving objects within a multi-channel, frame-like format. An example of this multi-stack event representation is illustrated in Figure 4. Based on this representation, our event encoder consists of a series of convolutional layers with stride to downsample the input event representation into the event latent $E_{cond}$, matching the dimension of the image latent $I_{cond}$.
@@ 3.4 Per-tile Denoising and Fusion
Due to the substantial memory demands of training and generating high-resolution video, most existing video diffusion models, such as Stable Video Diffusion (SVD) [7], are Latent Diffusion Models, meaning they operate in the latent space with downsampled spatial resolution rather than in the original pixel space. As illustrated in Figure 3, for an input RGB image/video of shape [$B$, $F$, 3, $H$, $W$], where $B$ is the batch size, $F$ is the number of frames, $W$ is the width, and $H$ is the height, a latent encoder ($\mathcal{E}$) and decoder ($\mathcal{D}$), typically based on a Variational Autoencoder (VAE), convert the input into image latents of shape [$B$, $F$, $C_{latent}$, $\frac{H}{d}$, $\frac{W}{d}$] and decode it back. Here, $d$ is the downsample factor, set to 8 in SVD, and $C_{latent}$ is the number of latent channels, set to 4 in SVD.
@@
Time Reversal [10] Pre-trained Video Diffusion 17.86 0.59 0.27 17.92 0.53 0.18 13.71 0.70 0.44
RIFE [15] Frame-only Video Frame Interpolation 23.30 0.83 0.09 25.34 0.76 0.05 16.93 0.76 0.35
CBMNet-Large [18] Event-based Video Frame Interpolation 26.24 0.79 0.22 28.73 0.84 0.09 22.26 0.86 0.39
RE-VDM (Ours) Adapted Video Diffusion 27.74 0.88 0.12 29.04 0.89 0.06 22.94 0.88 0.37
Table 2: We compare our method on real-world EVFI datasets against representative and SOTA methods, including the pre-trained video diffusion model (Time Reversal), frame-only method (RIFE), and event-based video frame interpolation method (CBMNet-Large). For CBMNet-Large, we use publicly available model checkpoints trained on the same BS-ERGB dataset as our approach. For the BS-ERGB and HQF datasets, we use a skip number of 3 as in the EVFI papers, we evaluate all interpolated frames. For our Clear-Motion test sequences, we use a skip number of 15 for evaluation on large motion scenarios, we evaluate the 4th, 8th, and 12th interpolated frames. To control for VAE encoding/decoding loss, as discussed in Section 4.2, we apply VAE encoding/decoding to all model outputs.
@@ 4 Experimental Results
4.1 Datasets and Training Details
To robustly evaluate the generalized frame interpolation performance of our approach, we select several real-world event camera datasets. The BS-ERGB dataset [37], the most cited for EVFI, provides aligned RGB and event video with large motion between frames, mainly featuring static cameras capturing dynamic objects. The HQF dataset [35] includes grayscale video aligned with events, with more sequences involving a moving event camera.
In addition to existing datasets, we include our self-collected Clear-Motion test sequences, which provide aligned RGB frames and events, capturing previously unseen large, distinct object motions between frames. This addition expands the range of scenarios covered beyond existing datasets, with details provided in the Supplementary Material.
The pre-trained video diffusion model we used is Stable Video Diffusion [7] for 14-frame image-to-video generation. We trained our model with an effective batch size of 64, using a batch size of 4 per GPU and a gradient accumulation factor of 16. Training was conducted solely on the BS-ERGB dataset, and the model was tested on other unseen datasets without fine-tuning. All training was performed on 4 NVIDIA RTX A6000 GPUs, each with 50GB of memory.
@@ 4.2 Evaluation Strategy
To evaluate generalized frame interpolation performance, we compare our approach with representative methods from three categories: frame-only interpolation (RIFE [15]), test-time optimization of pre-trained diffusion models (Time Reversal [10]), and event-based interpolation (CBMNet [18]). We use publicly available CBMNet checkpoints trained on the BS-ERGB dataset for evaluation. For Per-tile Denoising and Fusion in our method, which uses upsampled tiles to reduce detail loss and enhance event control accuracy, we apply an upsample factor of 2 for the BS-ERGB and Clear-Motion datasets and 3 for the HQF dataset.
For evaluation metrics, we use average frame PSNR [12], SSIM [39], and LPIPS [41] for the interpolated frames. Since SVD is a latent diffusion model that uses a VAE to encode and decode between latent and pixel space, we apply VAE encoding and decoding to all model outputs. This eliminates the impact of discrepancies between the VAE-decoded frame distribution and the ground truth frame distribution, accounting for differences in noise patterns, small details, and tone mapping.
@@ 4.3 Evaluation on Video Frame Interpolation
In this section, we present both qualitative and quantitative results that demonstrate the generalized performance of our proposed method compared to representative and SOTA methods for video frame interpolation, as shown in Table 2. As seen in the table, both Time Reversal and RIFE struggle with real-world video frame interpolation, as these methods do not utilize events between frames. Without event data, large or irregular motion between frames introduces too many degrees of freedom, making the interpolation problem ill-posed. In contrast, the SOTA event-based interpolation method, CBMNet-Large, can handle large motions by leveraging cues from events, similar to our approach. However, its performance on both the test set from the training dataset and unseen datasets is consistently lower than ours across all metrics. This difference is due to two factors: 1) as discussed in Section 1 and Table 1, the pre-trained video diffusion foundation model in our approach offers a superior data prior from its extensive training on large video generation datasets, and 2) our model’s backbone, a highly representative pre-trained video diffusion model, provides advantages over custom models typically used in EVFI.
The effectiveness of our approach on real-world, unseen data for video frame interpolation is clearly demonstrated in the qualitative comparisons in Figures 7, 8, 9, 10 and 11. Due to the ill-posed nature of interpolating frames with large motion, frame-only methods like RIFE and Time Reversal struggle to accurately reconstruct object locations and appearances. Notably, although Time Reversal’s reconstructions are not ideal, they show relatively higher temporal consistency compared to RIFE and CBMNet-Large. This highlights the strong data prior and temporal consistency advantage of pre-trained video diffusion foundation models, as discussed in Table 1. This observation supports our motivation to adapt these models for the EVFI task.
@@
While CBMNet-Large uses event data to guide interpolation and achieves similar object and texture placements, it lacks generalization on unseen object and camera motion compared to our method, as shown in Figures 7, 8, 9, 10, despite being trained on the same dataset. Additionally, CBMNet-Large suffers from frame-to-frame inconsistencies due to its frame-by-frame inference design, evident in Figure 11. CBMNet-Large produces frames with inconsistent textures and shapes on the ball and hand, whereas our method accurately reconstructs the challenging, regular shapes of both. This advantage is due to the strong data priors and foundational design of the pre-trained video diffusion model, allowing for consistent reconstruction across frames. Unlike CBMNet-Large, video diffusion models process multiple frames cohesively. These results underscore the effectiveness of our adaptation for generalized EVFI on real-world, unseen data and highlight the advantage of video diffusion models in achieving consistency across video frames.
Figure 7: An illustration showcasing the qualitative comparison on the Clear-Motion sequence Paper_Waving, which involves waving a paper with simple texture, with 11 skips between the start and end frames. We present the interpolated 4th, 7th, and 10th frames. (Zoom in for the best viewing experience)

### v2

@@ Abstract
Figure 1: We compare our proposed approach RE-VDM, which adapts a pre-trained video diffusion model for event-based video frame interpolation on unseen real-world data, with three baselines: frame-only interpolation method GIMM-VFI-R-P [14], test-time optimization of a pre-trained video diffusion model via Time Reversal [11], and event-based interpolation method CBMNet-Large [21], trained on the same dataset as our method. The left-most column shows the start frame, end frame, and reference interpolation frames overlaid with events. Leveraging data priors in the pre-trained video diffusion model and the diffusion process, along with motion guidance controlled by input events, our approach demonstrates superior generalization performance on unseen real-world frames with substantial motion.
@@ 2.1 Event Camera and Video Frame Interpolation
Event cameras are neuromorphic sensors that capture the brightness changes in the scene, and the brightness changes are primarily due to the camera’s motion and object motion in the scene. For this reason, events captured by event cameras can be used to compute motion-related information, for example, optical flow, motion segmentation, and ego-motion [6, 4, 3, 5, 29, 15, 37, 55, 54, 47]. Various methods have been proposed to reconstruct intensity videos from event streams alone, demonstrating that while feasible, the quality of such reconstructions is significantly limited by camera and object motion [9, 33]. Combining events with images extends beyond event-only applications, offering powerful capabilities for tasks like image deblurring and video frame interpolation/extrapolation [41, 42, 51, 10, 40]. This approach leverages events as cues to bridge the gap between frames, thus avoiding the ill-posed challenges faced by frame-only methods [18, 31, 25, 20, 1, 22, 45, 53, 48, 14] when interpolating between frames with substantial motion.
@@ 2.2 Diffusion and Video Diffusion Models
Here, $\alpha(t),\sigma(t)$ represents a noise schedule. Diffusion models have demonstrated remarkable image generation quality and diversity. Stable Diffusion [34] shifts the generation process to a low-dimensional latent space, greatly reducing computational cost. Trained on massive online text-image pair datasets [36], Stable Diffusion exhibits strong image generation capabilities. Video diffusion models [17, 8] add temporal layers on top of existing image diffusion models to jointly denoise multiple consistent image frames. Stable Video Diffusion (SVD) [8] is built on top of latent diffusion model and trained on massive video datasets, serving as a strong foundational model for video generation tasks.
@@ 2.3 Controllable Diffusion Generation Through Fine-Tuning
There have been numerous works on generation conditioned on additional user inputs or control [34, 35, 49, 46]. Stable Diffusion [34] uses cross-attention layers in the diffusion U-Net to inject text control. Dreambooth [35] showed that a diffusion model can be fine-tuned for conditional control. ControlNet [49] modifies a trainable copy of the original diffusion U-Net for conditional control. In addition to fine-tuning diffusion models for added controllability, various training-free methods have been proposed to alter the sampling process of image and video diffusion models for more flexible and generalized content synthesis [2, 11, 27, 52]. Latent diffusion models have limitations on generating a limited set of resolutions in a downsampled latent space. MultiDiffusion [2] provides a solution for resolution-agnostic generation by tile-based denoising. When combined with conditional control, MultiDiffusion [2] has shown to be effective in generating synthetic detailed images under the condition of fine-grained condition control [19, 43]. On the video generation side,  [11] proposes a test-time sampling method that empowers video diffusion models to control camera and object motion.  [52] enables consistent long-frame generation outside of the trained frame length of a video diffusion model.
@@ 3.1 Pipeline Overview
Figure 2 presents an overview of our pipeline RE-VDM, for adapting pre-trained video diffusion models to event-based video generation and interpolation tasks. The key difference is that interpolation applies our EVDS on both the start frame $I_{s}$ for forward-time video denoising/generation and the end frame $I_{e}$ for backward-time video denosing/generation. At each denoising step, forward-backward consistency is enforced to achieve accurate interpolation.
@@ 3.3 Events as the Control Condition
To introduce events as the control condition for the video generation process, we incorporate two main components: 1) Event Representation and 2) Event Encoder. For event representation, we adopt a multi-stack approach, inspired by [30, 26], which captures both fast- and slow-moving objects within a multi-channel, frame-like format. An example of this multi-stack event representation is illustrated in Figure 4. Based on this representation, our event encoder consists of a series of convolutional layers with stride to downsample the input event representation into the event latent $E_{cond}$, matching the dimension of the image latent $I_{cond}$.
@@ 3.4 Per-tile Denoising and Fusion
Due to the substantial memory demands of training and generating high-resolution video, most existing video diffusion models, such as Stable Video Diffusion (SVD) [7], are Latent Diffusion Models, meaning they operate in the latent space with downsampled spatial resolution rather than in the original pixel space. As illustrated in Figure 3, for an input RGB image/video of shape [$B$, $F$, 3, $H$, $W$], where $B$ is the batch size, $F$ is the number of frames, $W$ is the width, and $H$ is the height, a latent encoder ($\mathcal{E}$) and decoder ($\mathcal{D}$), typically based on a Variational Autoencoder (VAE), convert the input into image latents of shape [$B$, $F$, $C_{latent}$, $\frac{H}{d}$, $\frac{W}{d}$] and decode it back.
@@
Time Reversal [11] (ECCV’24) Video Diffusion 17.86 0.59 0.27 17.92 0.53 0.18 13.71 0.70 0.44
DynamiCrafter [46] (ECCV’24) Video Diffusion 15.47 0.54 0.40 15.99 0.56 0.35 11.42 0.62 0.52
RIFE [18] (ECCV’22) Frame-only VFI 23.30 0.83 0.09 25.34 0.76 0.05 16.93 0.76 0.35
PerVFI [45] (CVPR’24) Frame-only VFI 23.53 0.79 0.11 27.13 0.81 0.04 16.74 0.75 0.30
InterpAny-Clearer [53] (ECCV’24) Frame-only VFI 25.51 0.85 0.14 27.93 0.83 0.06 17.21 0.77 0.32
EMA-VFI [48] (CVPR’23) Frame-only VFI 26.01 0.86 0.18 28.31 0.84 0.10 17.79 0.78 0.44
GIMM-VFI-R-P [14] (NeurIPS’24) Frame-only VFI 25.56 0.84 0.08 27.96 0.83 0.04 17.77 0.77 0.26
CBMNet-Large [21] (CVPR’23) EVFI 26.24 0.79 0.22 28.73 0.84 0.09 22.26 0.86 0.39
RE-VDM (Ours) EVFI + Video Diffusion 27.74 0.88 0.12 29.04 0.89 0.06 22.94 0.88 0.37
Table 2: We compare our method on real-world EVFI datasets against representative and SOTA methods, including the pre-trained video diffusion model (Video Diffusion), frame-only methods (Frame-only VFI), and event-based video frame interpolation method (EVFI), where red indicates the best metric and blue indicates the second-best. For CBMNet-Large, we use publicly available model checkpoints trained on the same BS-ERGB dataset as our approach. For the BS-ERGB and HQF datasets, we use a skip number of 3 as in the EVFI papers, we evaluate all interpolated frames. For our Clear-Motion test sequences, we use a skip number of 15 for evaluation on large motion scenarios, we evaluate the 4th, 8th, and 12th interpolated frames. To control for VAE encoding/decoding loss, as discussed in Section 4.2, we apply VAE encoding/decoding to all model outputs.
@@ 4 Experimental Results
4.1 Datasets and Implementation Details
To robustly evaluate the generalized frame interpolation performance of our approach, we select several real-world event camera datasets. The BS-ERGB dataset [41], the most cited for EVFI, provides aligned RGB and event video with large motion between frames, mainly featuring static cameras capturing dynamic objects. The HQF dataset [39] includes grayscale video aligned with events, with more sequences involving a moving event camera. In addition to existing datasets, we include our self-collected Clear-Motion test sequences, capturing both object and camera motion. Detailed information is provided in the Supplementary Material. Clear-Motion is a self-collected dataset used solely for testing, eliminating the possibility of data leakage. This makes it ideal for robustly evaluating the zero-shot generalization performance of all models. The pre-trained video diffusion model we used is Stable Video Diffusion [7] for 14-frame image-to-video generation. We provide detailed training information in the supplementary material.
@@ 4.2 Evaluation Strategy
To evaluate generalized frame interpolation performance, we compare our approach with representative methods from three categories: frame-only interpolation, pre-trained video diffusion models, and event-based interpolation. We use publicly available CBMNet checkpoints trained on the BS-ERGB dataset for evaluation. For Per-tile Denoising and Fusion in our method, which uses upsampled tiles to reduce detail loss and enhance event control accuracy, we apply an upsample factor of 2 for the BS-ERGB and Clear-Motion datasets and 3 for the HQF dataset.
For evaluation metrics, we use average frame PSNR [13], SSIM [44], and LPIPS [50] for the interpolated frames. Since SVD is a latent diffusion model that uses a VAE to encode and decode between latent and pixel space, we apply VAE encoding and decoding to all model outputs. This eliminates the impact of discrepancies between the VAE-decoded frame distribution and the ground truth frame distribution, accounting for differences in small details and tone mapping.
@@ 4.3 Evaluation on Video Frame Interpolation
In this section, we present both quantitative and qualitative results demonstrating the effectiveness of our method compared to representative video frame interpolation approaches, as shown in Table 2. Our method outperforms all recent SOTA VFI baselines in terms of PSNR and SSIM. Qualitative comparisons in Figures 7 and 8 further show that recent frame-only VFI methods struggle with large and complex motion due to the ill-posed nature of the task. This limitation also affects pre-trained video diffusion models, which often exhibit a bias toward animation-style generation and suffer from spatial-temporal detail loss.
@@
In contrast, the SOTA event-based interpolation method, CBMNet-Large, can handle large motions by leveraging cues from events, similar to our approach. However, its performance on both the test set from the training dataset and unseen datasets is consistently lower than ours across all metrics. This difference is due to two factors: 1) as discussed in Section 1 and Table 1, the pre-trained video diffusion foundation model in our approach offers a superior data prior from its extensive training on large video generation datasets, and 2) our model’s backbone, a highly representative pre-trained video diffusion model, provides advantages over custom models typically used in EVFI.
@@
The effectiveness of our approach on real-world, unseen data for video frame interpolation is clearly demonstrated in the qualitative comparisons in Figure 7. Due to the ill-posed nature of interpolating frames with large motion, frame-only methods, despite often achieving better LPIPS scores, struggle to accurately reconstruct object locations and appearance. While CBMNet-Large uses event data to guide interpolation and achieves similar object and texture placements, it lacks generalization on unseen object and camera motion compared to our method, as shown in Figure 7, despite being trained on the same dataset. Additionally, CBMNet-Large suffers from frame-to-frame inconsistencies due to its frame-by-frame inference design, evident in Figure 8. CBMNet-Large produces frames with inconsistent textures and shapes on the ball and hand, whereas our method accurately reconstructs the challenging, regular shapes of both. This advantage is due to the strong data priors and foundational design of the pre-trained video diffusion model, allowing for consistent reconstruction across frames. These results support our unique contributions and motivation in adapting pre-trained video diffusion models for EVFI.
4.4 Ablation Study
In the Supplementary Material, we provide quantitative and qualitative ablation studies on two key components for adapting pre-trained VDMs to EVFI: 1) Comparing video generation and video interpolation using Two-Side Fusion. 2) Ablating the effect of upsampling in Per-tile Denoising and Fusion. Both quantitative and qualitative results demonstrate the effectiveness of the two main components in our approach for adapting pre-trained VDMs to the EVFI task.
Figure 7: The qualitative comparison on the Clear-Motion sequence Texture_Box, featuring translation and rotation of a textured box, demonstrates that our method achieves the most consistent and accurate frame reconstruction. Other methods experience significant detail loss, artifacts, and difficulty maintaining regular object shapes, particularly in the characters and the bottom-right coconut.
Figure 8: The qualitative comparison on the BS-ERGB test sequences, featuring fast motion of a hand and a rising, spinning ball, shows that our method is the only approach capable of maintaining the regular texture shapes on both the ball and hand in this challenging scenario.
5 Limitations
Despite the promising results, adapting pre-trained video diffusion models to the EVFI task presents several limitations. First, the pre-trained VAE encoder and decoder, used for conversion between pixel space and latent space, have limited representational power, often resulting in color shifts and loss of small details. Second, the adapted video diffusion model tends to smooth out certain regions, likely due to its difficulty in preserving fine textures within the downsampled latent space or the need for a prohibitively high upsampling factor to prevent smoothing. Finally, the diffusion-based methods are more memory-intensive and time-consuming than other methods, we provide the detailed model run time, memory comparison in the Supplementary Material. These limitations stem from the current state of pre-trained video diffusion foundation models. However, as the field of video generation evolves rapidly, we anticipate that many of these issues could be mitigated or resolved with newer, more advanced pre-trained models in the near future.
6 Conclusion
7 Acknowledgment
We greatly acknowledge NSF’s support under awards OISE 2020624 and BCS 2318255
Supplementary Material
8 Video Results
Please refer to our project page: https://vdm-evfi.github.io/ for video results, which clearly demonstrate that our reconstructions provide superior consistency and generalization compared to other baselines.
9 Video Generation Task Results
As explained in the main paper, our method supports Event-based Video Generation, an extrapolation task that relies on only one frame (start or end) and events, unlike interpolation, which uses both frames. This constraint in video generation leads to error accumulation in the generated video, as shown in the last video of the website. We present a comparison of video generation and interpolation results on the BS-ERGB dataset, for video generation, only the left-end frame next to skip frames and corresponding events are used to generate the skipped frames, as shown in Table 3, relying on information from only one side instead of both start and end frames causes the PSNR and SSIM metrics to drop significantly compared to the interpolation results.
Task BS-ERGB (3 skips)
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Video Generation 25.79 0.84 0.11
Video Interpolation 27.74 0.88 0.12
Table 3: Comparison of our pipeline’s performance on Video Generation and Video Interpolation tasks on the BS-ERGB dataset.
We also show a qualitative comparison in Figure 9, as we can see because the video generation task is a extrapolation task that only uses the information from one side (start or end frame) of frames and events, it cannot avoid hallucination for the occluded/missing parts that are not present in the condition image. In contrast, the video generation task can mitigate the hallucination well by using complementary information from both frames.
Figure 9: An example illustrating the difference between event-based video generation, which relies solely on the start frame, and event-based video interpolation, which uses both start and end frames to infer the interpolated frames. We present the 3rd, 6th, and 9th interpolated frames for the 11 skips interpolation between start and end frames on Clear-Motion test sequences. In the video generation scenario, the model hallucinates the occluded parts behind the paper due to the lack of information in the start frame. In contrast, video interpolation avoids hallucination as the end frame provides the necessary information.
In addition to hallucination, relying on frame information from only one side leads to error accumulation during generation, as shown in Figure 10, the results in video generation (incorrect color accumulation on the finger) are inconsistent with the information provided by the frame on the other side. In contrast, interpolation ensures the reconstructed video remains consistent with the information from both frames. The comparison of video consistency is best observed in the last video of the website, which compares event-based video generation with event-based video interpolation.
Figure 10: An example illustrating the difference in video consistency between event-based video generation and interpolation for 11 skips between start and end frames on Clear-Motion test sequences, we present the 5th, 8th, and 11th interpolated frames. In the video generation task, color errors on the finger accumulate over time, with the 11th frame (the final interpolated frame) failing to align with the information in the end frame. In contrast, interpolation reduces error accumulation and ensures consistency by leveraging information from both the start and end frames.
10 Clear-Motion Test Sequences
To robustly evaluate the zero-shot generalization performance of all models on unseen real-world event-based video frame interpolation scenarios, we collected the Clear-Motion Test Sequences solely for testing purposes.
10.1 Event-RGB Aligned Video Capture Setup
In this section, we will present the capture setup for our event-rgb aligned video sequences, as shown in Figure 11.
Figure 11: An illustration of our capture setup for the Clear-Motion test sequences of Event-RGB aligned video sequences. The setup consists of three main components: an event camera, an RGB camera, and a beam splitter to align the field of view for both cameras.
We use the Prophesee EVK4 HD as our event camera, offering a capture resolution of 1280 $\times$ 720. For the RGB camera, we use the BFS-U3-31S4C-C Blackfly S, which provides a resolution of 2048 $\times$ 1536 and supports up to 55 frames per second (fps). To align the field of view for both cameras, we utilize the Thorlabs CCM1-BS013 30 mm Cage Cube-Mounted Non-Polarizing Beam Splitter. Additionally, we perform spatial and temporal alignment to synchronize the events with the captured RGB frames. After the spatial alighment, our final captured RGB frames aligned with event are of resolution 940 $\times$ 720 and 40 fps.
10.2 Details of Data Sequence
The Clear-Motion test sequences are designed to include clear and large motions, encompassing both camera and object movements, as well as objects and motion patterns distinct from those found in most existing real-world Event-based Video Frame Interpolation (EVFI) datasets [41, 21, 39]. This setup enables a straightforward evaluation of the generalization and consistency of various video frame interpolation methods. Table 4 provides details of our collected test sequences, including explanations for each sequence. Category (i) represents sequences with object motion, while Category (ii) represents sequences with camera motion. We also include a Figure 12 to show some example data in our test sequences.
Figure 12: An illustration of example frames overlaid with events from Clear-Motion test sequences.
Sequence #Frames Explanation Category
Paper_Shifting 200 The translational and 3D rotational motion of a paper with a simple texture (i)
Paper_Waving 200 The waving and 3D rotational motion of a paper with a simple texture (i)
Paper_Deforming 200 The deformation motion of a paper with a simple texture (i)
Camera_Far 200 The moving cameras capturing distant objects (ii)
Camera_Close 200 The moving cameras capturing nearby objects (ii)
Checkerboard_Planar 200 The planar translation and rotation of a nearby dense checkerboard (i)
Checkerboard_Depth 200 The motion of a checkerboard along the depth direction (i)
Checkerboard_3D 200 The 3D translation and rotation of a nearby dense checkerboard (i)
Texture_Box 200 The translation and rotation of a nearby highly textured box (i)
Table 4: A detailed description of our collected Clear-Motion test sequences, with sequence name, number of frames and explanation of each sequence. Category (i) includes sequences with object motion, while Category (ii) includes sequences with camera motion.
11 The Impact of Input Upsampling
As discussed in the main paper, to mitigate the loss of appearance and motion control accuracy caused by the conversion between downsampled latent space and pixel space in Latent Diffusion Models (LDM) [7], we employ test-time optimization in the Per-tile Denoising and Fusion process. This involves first upsampling the input image and event representations by a specified factor and then breaking them into fixed-size overlapping tiles before feeding them into the video diffusion process. The performance comparison across different upsampling factors is shown in Table 5. As the upsampling factor increases from 1 to 2, our model’s performance improves significantly, with PSNR increasing by approximately 3 dB, SSIM by 0.11, and LPIPS decreasing by 0.03. These results demonstrate the effectiveness of upsampling in the Per-tile Denoising and Fusion process, enhancing both the details in reconstructed frames and the accuracy of event-based motion control.
Method BS-ERGB (3 skips)
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Ours_1 24.82 0.77 0.15
Ours_1.5 25.86 0.82 0.16
Ours_2 27.74 0.88 0.12
[…]
To qualitatively assess the effect of upsampling, Figure 13 shows that as the upsampling factor increases from 1 to 2, details on the human eyes and fingers (e.g., nails and textures) improve significantly. Both quantitative and qualitative results highlight the effectiveness of upsampling in the Per-tile Denoising and Fusion process, enhancing the realism and event-based motion control accuracy of video interpolation results.
Figure 13: An illustration showcasing the qualitative impact of input upsampling factors. The leftmost image is the reference frame. (a) shows the interpolated result with an upsampling factor of 1, (b) with a factor of 1.5, and (c) with a factor of 2. As the upsampling factor increases from 1 to 2, the details on the human eyes and fingers, highlighted by red arrows, improve significantly.
12 More Visual Results
In this section, we provide additional visual results showcasing qualitative comparisons between our method and the baselines, as shown in Figures 14, 15, 16, 17, 18, and 19.
13 Additional Implementation Details
In this section, we provide additional implementation details. The pre-trained video diffusion model we used is Stable Video Diffusion [7] for 14-frame image-to-video generation. We trained our model with an effective batch size of 64, using a batch size of 4 per GPU and a gradient accumulation factor of 16. Training was conducted solely on the BS-ERGB dataset, and the model was tested on other unseen datasets without fine-tuning. All training was performed on 4 NVIDIA RTX A6000 GPUs, each with 50GB of memory. For training, we use the AdamW optimizer [24] with a learning rate of $5\times 10^{-5}$ and parameters $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=1\times 10^{-8}$, and a weight decay of $1\times 10^{-2}$. Our model is trained on the BS-ERGB dataset for 72 hours to denoise noisy video latents for 3 and 11 skipped frames. All testing and inference on unseen data or datasets are conducted without fine-tuning, using the checkpoint trained on the BS-ERGB dataset.
During training and testing, the input to our adapted video diffusion model consists of 512 × 320 size tiles. Our model is trained solely to denoise/generate video latents using start frames and forward-time events. For testing and inference, we use an overlapping ratio of 0.1 for overlapping tiles and set the number of denoising steps in the video diffusion process to 25. The Per-tile Denoising and Fusion (for high-resolution frame reconstruction and event-based motion control) and Two-side Fusion (for converting video generation to interpolation) are both test-time optimization processes that do not require additional training.
14 Model Run Time, Memory, and Parameter Comparison
Table 6 reports testing results for all models run on a single NVIDIA RTX 4090 GPU. Each method generated $1024\times 576$ frames with run time averaged over 16 frames. VDM based methods (Time-Reversal, DynamiCrafter, and Ours) are more memory-intensive and time-consuming than other methods.
Method Run Time (s) Memory Usage (GB) #Parameters (M)
RIFE 0.5 0.9 9.8
CBMNet-Large 41.7 17.8 22.2
Time-Reversal 62.6 21.5 1524.6
PerVFI 9.3 5.2 13.9
InterpAny-Clearer 0.5 1.1 10.7
DynamiCrafter 92.0 18.3 1438.9
EMA-VFI 1.8 4.8 65.7
GIMM-VFI 3.0 9.5 19.8
Ours 200.1 17.2 2206.8
Table 6: Model Run Time, Memory, and Parameter comparison.
Figure 14: Additional baseline results on the Clear-Motion sequence Paper_Waving.
Figure 15: Additional baseline results on the BS-ERGB sequence as presented in the main paper.